The area under the normal curve between Z= -1.0 and Z = -2.0 is

A)  0.3413

B)  0.1359

C)  0.4772

D)  0.0228

2. Find Z if the area to the right of Z is 0.67

A)     0.9100

B)     0.0421

C)     -0.4500

D)     -0.390
Answer #1

solution

P( -1.0< Z < -2.0)

= P(Z < -2.0) - P(Z <-1.0 )

Using z table,  

=   0.1587-0.0228

area=0.1359

(B)

Using standard normal table,

P(Z > z) = 0.67

= 1 - P(Z < z) = 0.67

= P(Z < z) = 1 - 0.67

= P(Z < z ) =0.33

= P(Z < -0.4500 ) = 0.33  

z =-0.4500 ( using z table )